OpenStudy (anonymous):

What is the solution to the equations represented by these two lines? A. (4, 2) B. (2, 3) C. (2, 0) D. (0, 4)

OpenStudy (anonymous):

theres the graph

OpenStudy (anonymous):

@twinz1999 @imawesome101 @iceicebaby @undeadknight26 @UsArmy3947 @arabpride @highschoolmom2010 @nannaberry

OpenStudy (anonymous):

i have no clue chucky. srry :(

OpenStudy (anonymous):

is it c

undeadknight26 (undeadknight26):

srry man...

OpenStudy (anonymous):

thats ok nanna berry and thats ok undead knight

OpenStudy (anonymous):

i cant give you both a medal umm

OpenStudy (anonymous):

there a fan :) and a fan to nanna berry

OpenStudy (arabpride):

Ill brb and ill try to figure it out ok??? ill be right back in about 5-10 mins

OpenStudy (anonymous):

k

undeadknight26 (undeadknight26):

i think it is b because they intersect at 2,3.

undeadknight26 (undeadknight26):

hence it is the solution.

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 2y – x = 2 x – 2y = 8 A. consistent and independent B. coincident C. inconsistent D. not enough information

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

ima say d

undeadknight26 (undeadknight26):

no its c. the lines never meet.

OpenStudy (anonymous):

oops for got the graph thoe

OpenStudy (anonymous):

and they dont meet ok

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

sweet *gives undeadknight 10000000000 medals*

OpenStudy (anonymous):

i have a few more can u keep helpin

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 4y + x = 2 8y + 2x = 4 A. coincident B. consistent and independent C. inconsistent D. not enough information

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. 4y + x = 2 8y + 2x = 4 A. coincident B. consistent and independent C. inconsistent D. not enough information

OpenStudy (anonymous):

im gonna mention my friend

undeadknight26 (undeadknight26):

lol i kinda have a big test...here... https://www.desmos.com/calculator this should help u...

OpenStudy (anonymous):

@drewsheesley

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and i wanted u to see my kart

OpenStudy (anonymous):

hey undead how do u use the graph calculator

OpenStudy (anonymous):

drew or arab do you know it im tryin to get this done cause i have more

OpenStudy (anonymous):

alot more but drew and arab i can tc ina bit like 15 mins

OpenStudy (anonymous):

i thought it was b is that right

OpenStudy (arabpride):

OK im BACK!!!^.^

OpenStudy (anonymous):

yay :D

OpenStudy (arabpride):

ALright.......so u need help using the desmos graphing calculator???

OpenStudy (anonymous):

naa i dont get it si i need help on somthin else ima open a new ?

undeadknight26 (undeadknight26):

link me i might have a few secs to help.

OpenStudy (anonymous):

yo drew u have time for tc plz

OpenStudy (anonymous):

@drewsheesley